In a triangle ABC a:b:c=4:5:6.Find the ratio of the circumcircle to the incircle?

Grade:11

2 Answers

Vijay Mukati
8 years ago
Dear Student,

Since we know that, Area of triangle =abc/4R = rs,
From here R/r = 4sR2/abc,
Assume, a = 4x, b = 5x and c = 6x. Substitude this values in the above realation and Use the formula for R in terms of abc and find the ratio.

Thanks.
siva prasad reddy
8 years ago
/abc,Assume, a = 4x, b = 5x and c = 6x. Substitude this values in the above realation and Use the formula for R in terms of abc and find the ratio.Thanks. 2Since we know that, Area of triangle =abc/4R = rs,From here R/r = 4sR